OpenStudy (anonymous):

Proof \lim_{n \rightarrow ?\infty}n \sqrt{n ^{2}-n} =\frac{ 1 }{ 2 }

OpenStudy (anonymous):

\[\lim_{n \rightarrow ?\infty}n \sqrt{n ^{2}-n} =\frac{ 1 }{ 2 }\] So... what are your question?

OpenStudy (anonymous):

proof

hartnn (hartnn):

let y=1/n as n->infinity, y->0 replace all n by 1/y

OpenStudy (anonymous):

i don't understand

hartnn (hartnn):

check your question, its actually infinity http://www.wolframalpha.com/input/?i=lim+n-%3E+infinity++n+sqrt%28n%5E2-n%29&dataset=

OpenStudy (anonymous):

the question is: proof \lim_{n \rightarrow \infty}n \sqrt{n ^{2}-n} =\frac{ 1 }{ 2 }

OpenStudy (anonymous):

\[and another; proof \lim_{n \rightarrow \infty}0.99...9=1\]